OpenStudy (anonymous):

Please help me with this easy question. How do I find x and y from these 2 equations: 3x^2 + 3y^2 - 15 = 0 6xy + 3y^2 - 15 = 0 I'm having a brain fart right now.

OpenStudy (amistre64):

equate them

OpenStudy (amistre64):

3x^2 + 3y^2 - 15 = 6xy + 3y^2 - 15 3x^2 + 3y^2 = 6xy + 3y^2 3x^2 = 6xy 3x^2 - 6xy = 0 3x(x - 2y) = 0 when x = 0 or y = x/2

OpenStudy (amistre64):

3x^2 +0xy + 3y^2 - 15 0x^2 +6xy + 3y^2 - 15 1x^2 +0xy + 1y^2 - 5 0x^2 +6xy + 3y^2 - 15 1x^2 +0xy + 1y^2 - 5 0x^2 +1xy + 1/2y^2 - 5/2 x^2 = 5 - y^2 xy = (5-y^2)/2

OpenStudy (amistre64):

subbing in for x = sqrt(5-y^2) (24^2+1)y^2 - 10y + (25-5(24^2)) = 0 y = (10 +- sqrt(100-4(24^2+1)(25-5(24^2))))/(2(24^2+1)) y = abt 2.233 or -2.216 ------------------------------ y^2 = 5 - x^2 xsqrt(5-x^2) = -x^2/2 x^2(5-x^2) = -x^4/4 4x^2(5-x^2) = -x^4 20x^2 -4x^4 = -x^4 20x^2(20 -3x^2) = 0 x = 0, +-sqrt(20/3) maybe

OpenStudy (anonymous):

hmm, the solution says the points should be (2,1) , (-2,1) , (0, sqrt(5)) I think it should be something more obvious but I'm missing it..

OpenStudy (amistre64):

when x=0 is a fer sure 3(0)^2 + 3y^2 - 15 = 0 6(0)y + 3y^2 - 15 = 0 3y^2 = 15; y = +- sqrt(5)

OpenStudy (amistre64):

3x^2 + 3y^2 - 15 = 0 6xy + 3y^2 - 15 = 0 x = (15-3y^2)/6y might be a good sub

OpenStudy (anonymous):

so solve for x on the 2nd equaton and plug into 1st?

OpenStudy (amistre64):

yeah 3((15-3y^2)/6y)^2 + 3y^2 = 15 (15-3y^2)/6y)^2 + y^2 = 5 (15^2+9y^4-45y^2)/36y^2 + y^2 = 5 15^2+9y^4-45y^2 + 36y^4 = 5(36)y^2 5 - y^2 + y^4 = 4y^2 5 - 3y^2 + y^4 = 0

OpenStudy (amistre64):

solving the qudratic gets us a set: {-sqrt(5),-2,-1,1,2} to test out

OpenStudy (anonymous):

Ok thank you so much I will rework it

OpenStudy (amistre64):

good luck